OpenStudy (anonymous):

simplify radical square root of 200

OpenStudy (anonymous):

\[\sqrt{200} = \sqrt{100*2}\]

OpenStudy (anonymous):

\[\sqrt{200} = \sqrt{100}*\sqrt{2}\] \[\sqrt{200} = 10*\sqrt{2}\]

OpenStudy (anonymous):

thnxx